# A delayed interior area-to-height estimate for the Curve Shortening Flow

### v1

@@ 2 A local Harnack quantity
(2.33) $\partial_{t}\widetilde{\mathcal{H}}-\frac{1}{1+v^{2}}\partial_{x}^{2}\widetilde{\mathcal{H}}\geq 0$
within $J_{1}\times(0,\infty)$. From the Maximum Principle, we deduce the required ~$\widetilde{\mathcal{H}}(x,t)\geq 0$. ∎
@@
Recall that the conformal factor $u$ of a two–dimensional RF in local isothermal coordinates $(x,y)\in\mathbb{B}^{2}$ satisfies the Logarithmic Fast Diffusion Equation (LFDE)
(LFDE) $\partial_{t}u$ $=\Delta\log(u)$
where $\Delta=\partial_{x}^{2}+\partial_{y}^{2}$. Notice that (LFDE) is structurally similar to the equation satisfied by the spatial derivative $v$ of a GCSF,
(2.35) $\partial_{t}v$ $=\partial_{x}^{2}\!\left(\tan^{-1}(v)\!\right).$
@@
(2.37) $\partial_{t}\varphi$ $=\log\left(\Delta\varphi\right)+c$
@@
(2.39) $\partial_{t}\mathcal{A}$ $=\phi=\tan^{-1}(v)+\frac{\pi}{2}=\tan^{-1}\left(\partial_{x}^{2}\mathcal{A}\right)+\frac{\pi}{2}$
@@ 3 The delayed interior area-to-height estimate
Proof.
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times[0,T)\!)$ be a GCSF of the generality considered by Theorem 1.8. For any $10<\rho<$ 1, we consider the rescaled GCSF $u^{\rho}\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times[0,\rho^{-2}T)\!)$ given by
(3.1) $u^{\rho}(x,t)=\frac{1}{\rho}u(\rho x,\rho^{2}t),$
which starts from $u^{\rho}_{0}(x)=\rho^{-1}u_{0}(\rho x)$. Clearly then $u^{\rho}(t)\in\mathrm{C}^{0,1}(J_{1})$ for every $t\in[0,\rho^{-2}T)$ (it is, in fact, smooth up to the boundary), and we can readily compute that
(3.2) $\|u_{0}^{\rho}\|_{\mathrm{L}^{1}(J_{1})}$ $=\int_{x\in J_{1}}u^{\rho}_{0}(x)\,\mathrm{d}x=\frac{1}{\rho^{2}}\int_{x\in J_{\rho}}u_{0}(x)\,\mathrm{d}x=\frac{1}{\rho^{2}}\|u_{0}\|_{\mathrm{L}^{1}(J_{\rho})}$
@@
(3.3) $\left|u^{\rho}\left(0,\frac{t}{\rho^{2}}\right)\right|$ $=\frac{1}{\rho}|u(0,t)|.$
Let $\delta>0$. Then by applying the restricted Theorem 1.8 to the GCSF $u^{\rho}$, we have from (1.10) and (3.2)–(3.3) that
(3.4) $\frac{1}{\rho}|u(0,t)|\leq C+\frac{1}{2\rho^{2}}\|u_{0}\|_{\mathrm{L}^{1}(J_{\rho})}+\frac{\pi t}{2\rho^{2}}$
@@
(3.5) $\frac{t}{\rho^{2}}\geq(1+\delta)\cdot\frac{\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}}{\rho^{2}\pi}$
@@
If Theorem 1.8 holds for local GCSFs, then Theorem 1.8 also holds for GCSFs $u\in\mathrm{C}^{\infty}(J_{1}\times[0,T)\!)$ for which $u(t)\in\mathrm{C}^{0,1}(J_{1})$ for every $t\in[0,T)$.
Proof.
Let $u\in\mathrm{C}^{\infty}(J_{1}\times[0,T)\!)$ be a GCSF for which $u(t)\in\mathrm{C}^{0,1}(J_{1})$ for every $t\in[0,T)$, with $u_{0}:=u(0)$. We may suppose that $u_{0}\neq 0$ (in particular, that $\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}>0$), for otherwise the result is an immediate consequence of the Avoidance Principle (Theorem A.7). For any $\eta>0$, consider the Lipschitz function given by
(3.6) $\overline{u}_{0}(x)=|u_{0}(x)|+\eta>u_{0}(x)$
@@
(3.7) $u(x,t)\leq\overline{u}(x,t)$
for all $(x,t)\in J_{1}\times[0,T)$. Applying the restricted Theorem 1.8 to $\overline{u}$ with $\delta/2$, while noting that
(3.8) $\|\overline{u}_{0}\|_{\mathrm{L}^{1}(J_{1})}=\|{u}_{0}\|_{\mathrm{L}^{1}(J_{1})}+2\eta,$
@@
(3.9) $u(0,t)\leq\overline{u}(0,t)=|\overline{u}(0,t)|\leq C+\frac{1}{2}(\|{u}_{0}\|_{\mathrm{L}^{1}(J_{1})}+2\eta)+\frac{\pi t}{2}$
@@
(3.10) $t\geq(1+\delta)\cdot\frac{\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}}{\pi}\geq\left(1+\frac{\delta}{2}\right)\cdot\frac{\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}+2\eta}{\pi},$
as least for $\eta>0$ small enough. By taking $\eta\searrow 0$, we obtain the desired upper bound for $u(0,t)$. Repeating the argument above with $-u$ replacing $u$ yields the desired lower bound. ∎
The upshot of Claims 12 is, that to prove Theorem 1.8 in full generality, it suffices to prove the following version of the theorem.
@@
(3.12) $t_{\star}=\frac{\overline{\mathcal{A}}}{\pi}=\frac{\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}}{\pi}=\frac{1}{\pi}\int_{x\in J_{1}}u_{0}(x)\,\mathrm{d}x.$
Then, for all $\delta>0$, there exists a constant $C=C(\delta)<\infty$ such that the bound
(3.13) $|u(0,t)|\leq C+\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2}$
@@
(3.15) $\tan^{-1}(\partial_{x}u(y,t)\!)\leq\frac{\mathcal{A}(y,t)}{2t}.$
@@
(3.16) $\mathcal{A}(c_{t},t)=\frac{1}{2}\mathcal{A}(1,t)=\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2},$
using (2.20). By replacing $u(x,t)$ with $u(-x,t)$ if necessary, we may suppose that $c_{t}\geq 0$; consequently, for $y\leq 0\leq c_{t}$ we have that
(3.17) $\mathcal{A}(y,t)\leq\mathcal{A}(c_{t},t)\leq\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2}.$
@@
(3.18) $\tan^{-1}(\partial_{x}u(y,t)\!)$ $\leq\frac{1}{2t}\left(\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2}\right)$
$=\frac{\overline{\mathcal{A}}}{4t}+\frac{\pi}{4}$
$\leq\frac{\pi}{4(1+\delta)}+\frac{\pi}{4}<\frac{\pi}{2}.$
@@
(3.19) $\partial_{x}u(y,t)\leq 2C=2C(\delta)=\tan\left(\frac{\pi}{4(1+\delta)}+\frac{\pi}{4}\right),$
@@
To deduce the height bound, we use the upper gradient bound (3.19) to argue that the region trapped under $\gamma_{u}(t)$ must contain a trapezium, and then use the upper area bound (3.17) to control the area of the trapezium—see Fig. 3.1. Let $t\geq(1+\delta)\cdot t_{\star}$ be as before, and set $h=u(0,t)$. The upper gradient bound (3.19) implies that the line of gradient $2C$ through $(0,h)=(0,u(0,t)\!)\in\mathbb{R}^{2}$ must lie below $\gamma_{u}(t)$ to the left of the origin; in other words,
(3.20) $u(x,t)\geq h+2Cx$
for all $x\in(-1,0)$. Integrating (3.20) over $x\in(-1,0)$ and using the area bound (3.17), we get that
(3.21) $h-C\leq\int_{x\in(-1,0)}u(x,t)\,\mathrm{d}x=\mathcal{A}(0,t)\leq\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2},$
@@
(3.22) $|u(0,t)|=u(0,t)=h\leq C+\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2}.\qed$
[…]
Figure 3.1: The shaded region trapped below $\gamma_{u}(t)$ must contain the hatched trapezium, and so in particular the area of the hatched trapezium must be bounded from above by the area $\mathcal{A}(c_{t},t)=\mathcal{A}(1,t)/2=\overline{\mathcal{A}}/2+\pi t/2$ of the shaded region.
@@
(3.24) $C(\delta)$ $=\frac{1}{2}\tan\left(\frac{\pi}{2}-\frac{\pi\delta}{4(1+\delta)}\right)$
$=\frac{1}{2}\cot\left(\frac{\pi\delta}{4(1+\delta)}\right)$
$\leq\frac{1}{2}\csc\left(\frac{\pi\delta}{4(1+\delta)}\right).$
Using the lower bound $\sin(x)\geq 2x/\pi$ for $20<x<\pi/$ 2, we get the control
(3.25) $C(\delta)\leq\frac{1+\delta}{\delta}.$
By setting
(3.26) $t=(1+\delta_{t})\cdot t_{\star}\iff\delta_{t}=\frac{t-t_{\star}}{t_{\star}},$
@@
(3.27) $C(\delta_{t})\leq\frac{t}{t-t_{\star}}.$
@@
(3.29) $\|u(t)\|_{\mathrm{L}^{\infty}(\mathbb{R})}\leq\sqrt{t}\cdot\sqrt{\frac{2\|u_{0}\|_{\mathrm{L}^{1}(\mathbb{R})}}{t-t_{\star}}}.$
@@ 4 The instantaneous $\mathrm{L}^{p>1}_{\mathrm{loc}}$-to-$\mathrm{L}^{\infty}_{\mathrm{loc}}$–estimate
In this section, we shall use our delayed interior area-to-height estimate to deduce the instantaneous $\mathrm{L}^{p>1}_{\mathrm{loc}}$-to-$\mathrm{L}^{\infty}_{\mathrm{loc}}$–estimate of Theorem 1.11. The proof is essentially a transcription of [TY17, Theorem 4.1].
@@
As in Claim 1 in the proof of Theorem 1.8, it suffices to assume that $u(t)\in\mathrm{C}^{0,1}(J_{1})$ for each $t\in[0,T)$; we omit the details. Fix $t\in(0,T)$ and set $a=\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}=\pi\cdot t_{\star}$ for $t_{\star}:=\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}/\pi$. If $t\geq a$, then by Theorem 1.8 applied with $1+\delta^{\prime}=\pi$, we have that
(4.1) $|u(0,t)|\leq C(\delta^{\prime})+\frac{1}{2}a+\frac{\pi t}{2}\leq C(1+t)\leq C\left(1+\frac{\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p/(p-1)}}{t^{1/(p-1)}}+t\right).$
If $t<a$, then let $k\in\mathbb{R}_{>0}$ be such that
(4.2) $t=\|\overline{u}_{0}\|_{\mathrm{L}^{1}(J_{1})}=\int_{x\in J_{1}}\overline{u}_{0}(x)\,\mathrm{d}x=\int_{x\in J_{1}}(|u_{0}(x)|-k)^{+}\,\mathrm{d}x.$
@@
(4.4) $|u(0,t)|\leq k+\overline{u}(0,t)\leq k+C({\delta^{\prime}})+\frac{1}{2}t+\frac{\pi t}{2}.$
@@
(4.5) $\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p}$ $=\int_{x\in J_{1}}|u_{0}(x)|^{p}\,\mathrm{d}x\geq\int_{x\in\{|u_{0}|\geq k\}}|u_{0}(x)|^{p}\,\mathrm{d}x$
$\geq k^{p-1}\int_{x\in\{|u_{0}|\geq k\}}|u_{0}(x)|\,\mathrm{d}x\geq k^{p-1}\int_{x\in\{|u_{0}|\geq k\}}(|u_{0}(x)|-k)\,\mathrm{d}x$
$=k^{p-1}\int_{x\in J_{1}}(|u_{0}(x)|-k)^{+}\,\mathrm{d}x=k^{p-1}t$
@@
(4.6) $k\leq\frac{\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p/(p-1)}}{t^{1/(p-1)}},$
which combined with (4.4) yields
(4.7) $|u(0,t)|\leq\frac{\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p/(p-1)}}{t^{1/(p-1)}}+C({\delta}^{\prime})+\frac{1}{2}t+\frac{\pi t}{2}\leq C\left(1+\frac{\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p/(p-1)}}{t^{1/(p-1)}}+t\right).\qed$

### v2

@@ 2 A local Harnack quantity
$\partial_{t}\widetilde{\mathcal{H}}-\frac{1}{1+v^{2}}\partial_{x}^{2}\widetilde{\mathcal{H}}\geq 0$ (2.33)
within $J_{1}\times(0,\infty)$. From the Maximum Principle, we deduce the required $\widetilde{\mathcal{H}}(x,t)\geq 0$. ∎
@@
Recall that the conformal factor $u$ of a two–dimensional RF in local isothermal coordinates $(x,y)\in\mathbb{B}^{2}$ satisfies the Logarithmic Fast Diffusion Equation
$\partial_{t}u=\Delta\log(u),$ (LFDE)
where $\Delta:=\partial_{x}^{2}+\partial_{y}^{2}$. Notice that (LFDE) is structurally similar to the equation satisfied by the spatial derivative $v$ of a GCSF,
$\partial_{t}v=\partial_{x}^{2}[\tan^{-1}(v)].$ (2.35)
@@
$\partial_{t}\varphi=\log(\Delta\varphi)+c$ (2.37)
@@
$\partial_{t}\mathcal{A}=\phi=\tan^{-1}(v)+\frac{\pi}{2}=\tan^{-1}(\partial_{x}^{2}\mathcal{A})+\frac{\pi}{2}$ (2.39)
@@ 3 The delayed interior area-to-height estimate
?proofname?.
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times[0,T)\!)$ be a GCSF of the generality considered by Theorem 1.8. For any $0<\rho<1$, we consider the rescaled GCSF $u^{\rho}\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times[0,\rho^{-2}T)\!)$ given by
$u^{\rho}(x,t):=\frac{1}{\rho}u(\rho x,\rho^{2}t),$ (3.1)
which starts from $u^{\rho}_{0}(x):=\rho^{-1}u_{0}(\rho x)$. Clearly then $u^{\rho}(t)\in\mathrm{C}^{0,1}(J_{1})$ for every $t\in[0,\rho^{-2}T)$ (it is, in fact, smooth up to the boundary), and we can readily compute that
$\|u_{0}^{\rho}\|_{\mathrm{L}^{1}(J_{1})}=\int_{x\in J_{1}}u^{\rho}_{0}(x)\,\mathrm{d}x=\frac{1}{\rho^{2}}\int_{x\in J_{\rho}}u_{0}(x)\,\mathrm{d}x=\frac{1}{\rho^{2}}\|u_{0}\|_{\mathrm{L}^{1}(J_{\rho})}$ (3.2)
@@
$\left|u^{\rho}\left(0,\frac{t}{\rho^{2}}\right)\right|=\frac{1}{\rho}|u(0,t)|.$ (3.3)
Let $\delta>0$. Then by applying Theorem 1.8 to the GCSF $u^{\rho}$, we have from (1.10) and (3.2)–(3.3) that
$\frac{1}{\rho}|u(0,t)|\leq C+\frac{1}{2\rho^{2}}\|u_{0}\|_{\mathrm{L}^{1}(J_{\rho})}+\frac{\pi t}{2\rho^{2}}$ (3.4)
@@
$\frac{t}{\rho^{2}}\geq(1+\delta)\cdot\frac{\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}}{\rho^{2}\pi}\geq(1+\delta)\cdot\frac{\|u_{0}\|_{\mathrm{L}^{1}(J_{\rho})}}{\rho^{2}\pi},$ (3.5)
@@
If Theorem 1.8 holds for local GCSFs, then Theorem 1.8 also holds for GCSFs $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times[0,T)\!)$ for which $u(t)\in\mathrm{C}^{0,1}(J_{1})$ for every $t\in[0,T)$.
?proofname?.
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times[0,T)\!)$ be a GCSF for which $u(t)\in\mathrm{C}^{0,1}(J_{1})$ for every $t\in[0,T)$, with $u_{0}:=u(0)$. We may suppose that $u_{0}\neq 0$ (in particular, that $\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}>0$), for otherwise the result is an immediate consequence of the Avoidance Principle (Theorem A.7). For any $\eta>0$, consider the positive Lipschitz function given by
$\overline{u}_{0}(x):=|u_{0}(x)|+\eta>u_{0}(x)$ (3.6)
@@
$u(x,t)\leq\overline{u}(x,t)$ (3.7)
for all $(x,t)\in J_{1}\times[0,T)$. Applying Theorem 1.8 to $\overline{u}$ with $\delta/2$, while noting that
$\|\overline{u}_{0}\|_{\mathrm{L}^{1}(J_{1})}=\|{u}_{0}\|_{\mathrm{L}^{1}(J_{1})}+2\eta>\|{u}_{0}\|_{\mathrm{L}^{1}(J_{1})}>0,$ (3.8)
@@
$u(0,t)\leq\overline{u}(0,t)=|\overline{u}(0,t)|\leq C+\frac{1}{2}(\|{u}_{0}\|_{\mathrm{L}^{1}(J_{1})}+2\eta)+\frac{\pi t}{2}$ (3.9)
@@
$t\geq(1+\delta)\cdot\frac{\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}}{\pi}\geq\left(1+\frac{\delta}{2}\right)\cdot\frac{\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}+2\eta}{\pi},$ (3.10)
as least for $\eta>0$ small enough. By taking $\eta\searrow 0$, we obtain the desired upper bound for $u(0,t)$. Repeating the argument above with u-u replacing $u$ yields the desired lower bound. ∎
The upshot of Claims 12 is, that to prove Theorem 1.8 in full generality, it suffices to prove the following version of the theorem:
@@
$t_{\star}:=\frac{\overline{\mathcal{A}}}{\pi}=\frac{\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}}{\pi}=\frac{1}{\pi}\int_{x\in J_{1}}u_{0}(x)\,\mathrm{d}x.$ (3.12)
Then for all $\delta>0$, there exists a constant $C=C(\delta)<\infty$ such that the bound
$|u(0,t)|\leq C+\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2}$ (3.13)
@@
$\tan^{-1}[\partial_{x}u(y,t)]\leq\frac{\mathcal{A}(y,t)}{2t}.$ (3.15)
@@
$\mathcal{A}(c_{t},t)=\frac{1}{2}\mathcal{A}(1,t)=\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2},$ (3.16)
using (2.20). By replacing $u(x,t)$ with u(x,t)u(-x,t) if necessary, we may suppose that $c_{t}\geq 0$; consequently, for $y\leq 0\leq c_{t}$ we have that
$\mathcal{A}(y,t)\leq\mathcal{A}(c_{t},t)=\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2}.$ (3.17)
@@
$\tan^{-1}[\partial_{x}u(y,t)]$ $\leq\frac{1}{2t}\left(\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2}\right)=\frac{\overline{\mathcal{A}}}{4t}+\frac{\pi}{4}\leq\frac{\pi}{4(1+\delta)}+\frac{\pi}{4}<\frac{\pi}{2}.$ (3.18)
@@
$\partial_{x}u(y,t)\leq 2C=2C(\delta):=\tan\left[\frac{\pi}{4(1+\delta)}+\frac{\pi}{4}\right],$ (3.19)
@@
To deduce the height bound, we use the upper gradient bound (3.19) to argue that the region trapped under $\gamma(t):=\operatorname{graph}(u(t)\!)$ must contain a trapezium, and then use the upper area bound (3.17) to control the area of the trapezium—see Fig. 3.1. Let $t\geq(1+\delta)\cdot t_{\star}$ be as before, and set $h:=u(0,t)$. The upper gradient bound (3.19) implies that the line of gradient $2C$ through $(0,h)=(0,u(0,t)\!)\in\mathbb{R}^{2}$ must lie below $\gamma(t)$ to the left of the origin; in other words,
$u(x,t)\geq h+2Cx$ (3.20)
for all 1<x<0-1<x<0. Integrating (3.20) over $x\in(-1,0)$ and using the area bound (3.17), we get that
$h-C=\int_{x\in(-1,0)}[h+2Cx]\,\mathrm{d}x\leq\int_{x\in(-1,0)}u(x,t)\,\mathrm{d}x=\mathcal{A}(0,t)\leq\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2},$ (3.21)
@@
$|u(0,t)|=u(0,t)=h\leq C+\frac{\overline{\mathcal{A}}}{2}+\frac{\pi t}{2}.\qed$ (3.22)
@@
$C(\delta)=\frac{1}{2}\tan\left[\frac{\pi}{2}-\frac{\pi\delta}{4(1+\delta)}\right]=\frac{1}{2}\cot\left[\frac{\pi\delta}{4(1+\delta)}\right]\leq\frac{1}{2}\csc\left[\frac{\pi\delta}{4(1+\delta)}\right].$ (3.24)
Together with the lower bound $\sin(x)\geq 2x/\pi$ for $0<x<\pi/2$, we deduce that
$C(\delta)\leq\frac{1+\delta}{\delta}.$ (3.25)
Therefore, by choosing $\delta_{t}>0$ so that
$t=:(1+\delta_{t})\cdot t_{\star}\iff\delta_{t}:=\frac{t-t_{\star}}{t_{\star}},$ (3.26)
@@
$C(\delta_{t})\leq\frac{t}{t-t_{\star}}.$ (3.27)
@@
$\|u(t)\|_{\mathrm{L}^{\infty}(\mathbb{R})}\leq\sqrt{t}\cdot\sqrt{\frac{2\|u_{0}\|_{\mathrm{L}^{1}(\mathbb{R})}}{t-t_{\star}}}.$ (3.29)
[…]
?figurename? 3.1: The shaded region trapped below $\gamma(t):=\operatorname{graph}(u(t)\!)$ must contain the hatched trapezium, and so in particular the area of the hatched trapezium must be bounded from above by the area $\mathcal{A}(c_{t},t)=\mathcal{A}(1,t)/2=\overline{\mathcal{A}}/2+\pi t/2$ of the shaded region.
@@ 4 The instantaneous $\mathrm{L}^{p>1}_{\mathrm{loc}}$-to-$\mathrm{L}^{\infty}_{\mathrm{loc}}$–estimate
In this section, we shall use our delayed interior area-to-height estimate (Theorem 1.8) to deduce the instantaneous $\mathrm{L}^{p}_{\mathrm{loc}}$-to-$\mathrm{L}^{\infty}_{\mathrm{loc}}$–estimate of Theorem 1.12, where $1<p<\infty$. The proof is essentially a transcription of [TY17, Theorem 4.1].
@@
As in Claim 1 in the proof of Theorem 1.8, it suffices to assume that $u(t)\in\mathrm{C}^{0,1}(J_{1})$ for each $t\in[0,T)$; we omit the details. Fix $t\in(0,T)$ and set $a=\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}=\pi\cdot t_{\star}$ for $t_{\star}:=\|u_{0}\|_{\mathrm{L}^{1}(J_{1})}/\pi$.
If $t\geq a$, then by Theorem 1.8 with $1+\delta^{\prime}=\pi$ we have that
$|u(0,t)|\leq C(\delta^{\prime})+\frac{1}{2}a+\frac{\pi t}{2}\leq C(1+t)\leq C\left(1+\frac{\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p/(p-1)}}{t^{1/(p-1)}}+t\right).$ (4.1)
If $t<a$, then let $k\in\mathbb{R}_{>0}$ be such that the function $\overline{u}_{0}:=(|u_{0}|-k)^{+}$ satisfies
$\|\overline{u}_{0}\|_{\mathrm{L}^{1}(J_{1})}=\int_{x\in J_{1}}\overline{u}_{0}(x)\,\mathrm{d}x=\int_{x\in J_{1}}(|u_{0}(x)|-k)^{+}\,\mathrm{d}x=t.$ (4.2)
@@
$|u(0,t)|\leq k+\overline{u}(0,t)\leq k+C({\delta^{\prime}})+\frac{1}{2}t+\frac{\pi t}{2}.$ (4.4)
@@
$\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p}$ $=\int_{x\in J_{1}}|u_{0}(x)|^{p}\,\mathrm{d}x$ (4.5)
$\geq\int_{x\in\{|u_{0}|\geq k\}}|u_{0}(x)|^{p}\,\mathrm{d}x\geq k^{p-1}\int_{x\in\{|u_{0}|\geq k\}}|u_{0}(x)|\,\mathrm{d}x$
$\geq k^{p-1}\int_{x\in\{|u_{0}|\geq k\}}(|u_{0}(x)|-k)\,\mathrm{d}x=k^{p-1}\int_{x\in J_{1}}(|u_{0}(x)|-k)^{+}\,\mathrm{d}x=k^{p-1}t,$
@@
$k\leq\frac{\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p/(p-1)}}{t^{1/(p-1)}},$ (4.6)
which combined with (4.4) gives
$\begin{aligned} |u(0,t)|\leq\frac{\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p/(p-1)}}{t^{1/(p-1)}}+C({\delta}^{\prime})+\frac{1}{2}t+\frac{\pi t}{2}\leq C\left(1+\frac{\|u_{0}\|_{\mathrm{L}^{p}(J_{1})}^{p/(p-1)}}{t^{1/(p-1)}}+t\right).\end{aligned}\qed$ (4.7)